# Advancing Depth Anything Model for Unsupervised Monocular Depth Estimation in Endoscopy

### v1

@@ Abstract
Depth estimation is a cornerstone of 3D reconstruction and plays a vital role in minimally invasive endoscopic surgeries. However, most current depth estimation networks rely on traditional convolutional neural networks, which are limited in their ability to capture global information. Foundation models offer a promising avenue for enhancing depth estimation, but those currently available are primarily trained on natural images, leading to suboptimal performance when applied to endoscopic images. In this work, we introduce a novel fine-tuning strategy for the Depth Anything Model and integrate it with an intrinsic-based unsupervised monocular depth estimation framework. Our approach includes a low-rank adaptation technique based on random vectors, which improves the model’s adaptability to different scales. Additionally, we propose a residual block built on depthwise separable convolution to compensate for the transformer’s limited ability to capture high-frequency details, such as edges and textures. Our experimental results on the SCARED dataset show that our method achieves state-of-the-art performance while minimizing the number of trainable parameters. Applying this method in minimally invasive endoscopic surgery could significantly enhance both the precision and safety of these procedures.
@@ I INTRODUCTION
Depth estimation is a crucial component of augmented reality navigation systems in minimally invasive endoscopic surgery, where accurate 3D depth information is essential for enhancing surgical precision and safety [1, 2]. Given the difficulties in obtaining ground truth from endoscopic images and the fact that monocular endoscopes are more flexible and practical compared to stereo endoscopes, unsupervised monocular depth estimation (UMDE) algorithms have attracted broader research interest [3, 4]. These algorithms typically transform the depth estimation task into an image synthesis problem between adjacent viewpoints, using image synthesis error to guide network training [5]. Despite their potential, conventional depth estimation algorithms face significant challenges when applied to endoscopic images due to factors such as lighting variations, tissue deformation, and sparse textures [6].
@@
Recently, a foundation model for depth estimation in natural images (Depth Anything Model) has been proposed [12]. However, as shown in Fig. 1, this model exhibits significant performance degradation when directly applied to endoscopic images [13]. Furthermore, due to the scarcity of annotated endoscopic data, training a foundation model for endoscopic depth estimation from scratch is impractical. Therefore, fine-tuning the Depth Anything Model for specific endoscopic scenarios has become an effective approach to addressing this issue.
There are numerous strategies for fine-tuning foundation models, with Low-Rank Adaptation (LoRA) being one of the most commonly used methods [14]. However, LoRA does not always perform optimally and some variations show better performance in certain tasks. To address this, we explore a novel fine-tuning strategy aimed at enhancing performance without increasing the number of parameters in LoRA. Inspired by recent studies that highlight the surprising effectiveness of models using random weights and projections [15, 16], we propose a low-rank adaptation method based on random vectors (RVLoRA) that enhances model performance without adding additional trainable parameters. Additionally, transformers tend to focus more on low-frequency information, resulting in an insufficient ability to extract details and local features such as edges and textures [17, 18]. To address this, we propose a residual block based on depthwise separable convolution (Res-DSC) to achieve higher performance by enhancing locality.
@@
  • We designed a novel fine-tuning strategy, RVLoRA, for depth estimation foundation model. This strategy enhances the model’s adaptability to different scales by introducing random scaling vectors.
@@
  • Experiments conducted on the SCARED dataset demonstrate that our approach achieves state-of-the-art results with the minimal trainable parameters.
@@ II-A Unsupervised Monocular Depth Estimation
Despite the advances these methods have made for natural images, most of them rely on the assumption of photometric consistency, which renders them less effective in endoscopic scenarios with varying illumination. To address this, Ozyoruk et al. [20] and Shao et al. [7] first adjusted the lighting of images before performing depth estimation. Li et al. [6] improved the results by employing intrinsic decomposition to separate illumination variations and using reflectance consistency. Besides this line of research, Yang et al. [21] proposed a lightweight model combining CNNs and transformers, but this came at the cost of depth estimation accuracy. Cui et al. [13] combined foundation model techniques for depth estimation, achieving relatively better results.
@@ III METHOD
In this section, we first review the framework of image intrinsic-based unsupervised monocular depth estimation, which we have integrated with the Depth Anything Model. Following that, we provide a detailed explanation of our proposed fine-tuning strategy. Fig. 2 illustrates our network framework.
@@ III-A Image Intrinsic-Based UMDE
In classic UMDE framework [4], taking one frame as the target image $I_{t}$ and adjacent frames as source images $I_{s}$. A depth estimation network is used to estimate the depth information $D_{t}$ of the target image, and a pose estimation network is used to estimate the camera’s pose transformation matrix $T_{t\rightarrow s}$ between adjacent frames. Using known camera intrinsic matrix $K$, the position correspondence between the $I_{t}$ and $I_{s}$ is determined:
@@ III-B Advancing Depth Anything Model for UMDE
The depth estimation networks typically employed in classic UMDE methods are standard convolutional neural networks, which lack sufficient global perception capabilities for accurately capturing depth information. Fine-tuning a foundation model is a straightforward approach to enhance long-range dependencies without significantly increasing the number of trainable parameters. Although the Depth Anything Model [12] has been proposed for depth estimation, it was trained on natural images, and applying it directly to endoscopic images leads to poor performance. To address this, we propose a novel fine-tuning strategy specifically tailored for endoscopic images. Our approach introduces two key components: low-rank adaptation based on random vectors and residual blocks utilizing depthwise separable convolution.
@@ III-B1 Low-rank Adaptation Based on Random Vectors (RVLoRA)
Recent studies have shown that models utilizing random weights and predictions have remarkable effectiveness, and introducing random vectors can significantly enhance the model’s adaptive capability [15, 16]. Furthermore, various deep learning techniques rely on affine transformations of learned features. Freezing the randomly initialized parameters can enhance the network’s ability to capture these affine transformations [25]. Additionally, in monocular depth estimation tasks, an inherent issue is scale ambiguity. Introducing random vectors for scale adjustment can enhance the model’s ability to adapt to varying scales. Inspired by this, we propose a new parameter-efficient fine-tuning method: low-rank adaptation based on random vectors (RVLoRA).
@@
where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, $b\in\mathbb{R}^{m\times 1}$, $a\in\mathbb{R}^{r\times 1}$, with rank $r\ll\min(m,n)$. In this method, the low-rank matrices $B$ and $A$ are trainable, while the scaling vectors $b$ and $a$ are random and frozen, formally represented as diagonal matrices $\Lambda_{b}$ and $\Lambda_{a}$. This approach effectively performs random scaling on the rows of the low-rank matrices $B$ and $A$, thereby enhancing the model’s adaptability across layers.
@@ III-B2 Residual Block Based on Depthwise Separable Convolution (Res-DSC)
Transformers offer advantages like dynamic attention, global context awareness, and strong generalization capabilities. However, they often struggle to extract fine details and local features. On the other hand, convolutions excel at capturing local features and come with inherent benefits such as shift, scale, and distortion invariance, though they are limited in their ability to perceive global context. Therefore, combining CNNs with Transformers creates a complementary effect, enhancing overall model performance.
Inspired by this, we designed a residual block based on depthwise separable convolution (Res-DSC), as illustrated in Fig. 4. This structure consists of three convolutional layers: the first layer is a 1$\times$1 convolution to reduce the number of channels; the second layer is a 3$\times$3 depthwise separable convolution, which has fewer parameters, faster computation, and a more streamlined model compared to traditional convolution; the final layer is a 1$\times$1 convolution to restore the number of channels. To reduce the model’s parameters while ensuring the accuracy of depth estimation, we only add Res-DSC after the 3rd, 6th, 9th, and 12th transformer blocks. In Fig. 2, only the Res-DSC blocks after the 9th and 12th transformer blocks are shown, while the others are omitted for brevity.
@@ III-B3 Decoder
The decoder part adopts the same structure as Depth Anything Model. The features from the last four layers of the encoder are fed into the decoder. Through the frozen Reassemble &\&& Fusion module in the decoder, features at different resolutions are obtained and then aggregated. These aggregated features are then fed into the trainable Depth-Head module for the final dense prediction, resulting in depth estimation outputs at four scales.
@@ III-C Total Training Framework
Figure 5: The visual results of the comparative experiments. It is evident that our method performs better in capturing details such as object edges. In particular, the regions highlighted by the black boxes demonstrate a significant advantage of our approach.
@@
SfMLearner [4] CVPR’17 0.086 1.021 7.553 0.121 0.925 31.60 31.60
Monodepth2 [5] ICCV’19 0.066 0.577 5.781 0.093 0.961 14.84 14.84
Endo-SfMLearner [20] MedIA’21 0.068 0.679 6.227 0.098 0.955 14.84 14.84
AF-SfMLearner [7] MedIA’22 0.060 0.443 4.964 0.082 0.973 14.84 14.84
IID-SfMLearner [6] JBHI’24 0.058 0.435 4.820 0.080 0.969 14.84 14.84
Depth Anything [12] CVPR’24 0.084 0.847 6.711 0.110 0.930 97.50 97.50
EndoDAC [13] MICCAI’24 0.052 0.362 4.464 0.073 0.979 99.09 1.66
Ours - 0.048 0.315 4.172 0.068 0.982 98.80 1.38
@@ IV-B Dataset and Metrics
We conducted our experiments on the SCARED dataset, which was collected using the da Vinci Xi endoscope during the dissection of fresh pig abdomens [30]. High-quality depth ground truth was obtained using structured light encoding during the acquisition process. Following the splitting strategy from prior works [6, 7], we divided the SCARED dataset into 15,351 frames for training, 1,705 frames for validation, and 551 frames for testing.
We adopted commonly used evaluation metrics such as Abs Rel, Sq Rel, RMSE, RMSE log, and Threshold $\delta$, with their calculation methods shown in Table I. Similar to previous methods [6, 7], during the evaluation, we scale the predicted depth map via the median scaling, which can be expressed as:
@@
where $d$ represents the predicted depth, and $d^{*}$ represents the ground truth. The upper limit for scaling the depth map is set to 150 millimeters, which effectively covers almost all pixels in the SCARED dataset.
Figure 6: Visualization of pose estimation results. The blue line represents the ground truth, while the green line represents the predicted results. It can be seen that our results exhibit less trajectory drift and align more closely with the ground truth.
@@ IV-C Comparison Experiments
We compared our method with seven unsupervised depth estimation methods: SfmLearner [4], Monodepth2 [5], Endo-SfmLearner [20], AF-SfmLearner [7], IID-SfmLearner [6], Depth Anything [12], and EndoDAC [13]. Among these, EndoDAC is currently the state-of-the-art method on the SCARED dataset. The first three methods are reproduced from scratch following the code provided by the authors, and the last four methods are tested using the optimal model provided by the authors. Table II presents the results of the comparative experiments.
@@
Fig. 5 shows some representative results for the compared methods. Since Depth Anything Model is trained on large-scale natural images, it performs well on objects commonly found in natural scenes, such as instruments. However, overall, the depth maps estimated by our method exhibit clearer object edges and demonstrate superior performance in capturing details.
@@ IV-D Ablation Study
To further demonstrate the effectiveness of the proposed modules, we conducted ablation experiments, as shown in Table IV. The first row in the table represents the results obtained by directly testing the Depth Anything Model without any fine-tuning. The subsequent rows show the results of progressively adding the proposed two modules. The experimental results indicate that both modules effectively improve the depth estimation performance.
In our method, the Res-DSC modules are evenly distributed across the 12 transformer modules of the encoder. For instance, when using three Res-DSC modules, they are placed at the 4th, 8th, and 12th layers. Therefore, to determine the optimal number of Res-DSC modules, we conducted ablation experiments. The results, shown in Table V, indicate that using 4 Res-DSC modules achieves the best performance while maintaining a relatively low parameter count.
TABLE IV: The Ablation Study Results.
@@
TABLE V: The Ablation Study on The Number of Res-DSC Modules
@@
TABLE VI: The Results of Pose Estimation.
Method ATE (Seq 1$\downarrow$ ATE (Seq 2$\downarrow$
EndoDAC[13] 0.0779 0.0521
Ours 0.0771 0.0505

### v2

@@ Abstract
Depth estimation is a cornerstone of 3D reconstruction and plays a vital role in minimally invasive endoscopic surgeries. However, most current depth estimation networks rely on traditional convolutional neural networks, which are limited in their ability to capture global information. Foundation models offer a promising approach to enhance depth estimation, but those models currently available are primarily trained on natural images, leading to suboptimal performance when applied to endoscopic images. In this work, we introduce a novel fine-tuning strategy for the Depth Anything Model and integrate it with an intrinsic-based unsupervised monocular depth estimation framework. Our approach includes a low-rank adaptation technique based on random vectors, which improves the model’s adaptability to different scales. Additionally, we propose a residual block built on depthwise separable convolution to compensate for the transformer’s limited ability to capture local features. Our experimental results on the SCARED dataset and Hamlyn dataset show that our method achieves state-of-the-art performance while minimizing the number of trainable parameters. Applying this method in minimally invasive endoscopic surgery can enhance surgeons’ spatial awareness, thereby improving the precision and safety of the procedures.
@@ I INTRODUCTION
Depth estimation is a crucial component of augmented reality navigation systems in minimally invasive endoscopic surgery, where accurate 3D depth information is essential for enhancing surgical precision and safety [1, 2]. Given the difficulties in obtaining ground truth from endoscopic images and the fact that monocular endoscopes are more flexible and practical compared to stereo endoscopes, unsupervised monocular depth estimation (UMDE) algorithms have attracted broader research interest [3, 4]. These algorithms typically transform the depth estimation task into an image synthesis problem between adjacent viewpoints, using image synthesis error to guide network training [5]. However, conventional depth estimation algorithms face significant challenges when applied to endoscopic images due to factors such as lighting variations and sparse textures [6].
@@
Recently, a foundation model (Depth Anything Model) for depth estimation in natural images has been proposed [12]. However, as shown in Fig. 1, this model exhibits significant performance degradation when directly applied to endoscopic images [13]. Furthermore, due to the scarcity of annotated endoscopic data, training a foundation model for endoscopic depth estimation from scratch is impractical. Therefore, fine-tuning the Depth Anything Model for specific endoscopic scenarios has become an effective approach to addressing this issue.
There are numerous strategies for fine-tuning foundation models, with Low-Rank Adaptation (LoRA) being one of the most commonly used methods [14]. However, LoRA does not always perform optimally and some variations show better performance in certain tasks. To address this, we explore a novel fine-tuning strategy aimed at enhancing performance without increasing the number of parameters in LoRA. Inspired by recent studies that highlight the surprising effectiveness of models using random weights and projections [15, 16], we propose a low-rank adaptation method based on random vectors (RVLoRA) that enhances model performance without adding additional trainable parameters. Additionally, transformer models tend to focus more on low-frequency information, resulting in insufficient ability to extract texture details and local features [17, 18]. To address this, we propose a residual block based on depthwise separable convolution (Res-DSC) to achieve higher performance by enhancing locality.
@@
  • We designed a novel fine-tuning strategy, RVLoRA, for depth estimation foundation model. This strategy enhances the model’s adaptability to different scales by introducing random scaling vectors, without adding additional trainable parameters.
@@
  • Experiments conducted on SCARED dataset and Hamlyn dataset demonstrate that our approach achieves state-of-the-art results with the minimal trainable parameters.
@@ II-A Unsupervised Monocular Depth Estimation
Despite the advances these methods have made for natural images, most of them rely on the assumption of photometric consistency, which renders them less effective in endoscopic scenarios with varying illumination. To address this, Ozyoruk et al. [20] and Shao et al. [7] first adjusted the lighting of images before performing depth estimation. Li et al. [6] improved the results by employing intrinsic decomposition to separate illumination variations and using reflectance consistency. Besides this line of research, Batlle et al. [21] used photometric stereo technique to achieve depth estimation of the human colon, but this method only performed well on the cavitary endoscope dataset. Cui et al. [13] combined foundation model techniques for depth estimation, achieving relatively better results.
@@ III METHOD
In this section, we first review the framework of image intrinsic-based unsupervised monocular depth estimation, which serves as the baseline for our method. We then describe how we fine-tune the foundational model to obtain a better depth estimation. Fig. 2 illustrates our network framework.
@@ III-A Image Intrinsic-Based UMDE
In classic UMDE framework [4], taking one frame as the target image $I_{t}$ and adjacent frames as source images $I_{s}$, a depth estimation network is used to estimate the depth information $D_{t}$ of the target image, and a pose estimation network is used to estimate the camera’s pose transformation matrix $T_{t\rightarrow s}$ between adjacent frames. Using known camera intrinsic matrix $K$, the position correspondence between the $I_{t}$ and $I_{s}$ is determined:
@@ III-B Advancing Depth Anything Model for UMDE
The depth estimation networks typically employed in IID-SfMLearner and other classic UMDE methods are standard convolutional neural networks, which lack sufficient global perception capabilities for accurately capturing depth information. Fine-tuning a foundation model is a straightforward approach to enhance long-range dependencies without significantly increasing the number of trainable parameters. Although the Depth Anything Model [12] has been proposed for depth estimation, it was trained on natural images, and applying it directly to endoscopic images leads to poor performance. To address this, we propose a novel fine-tuning strategy specifically tailored for endoscopic images. Our approach introduces two key components: low-rank adaptation based on random vectors and residual blocks utilizing depthwise separable convolution.
@@ III-B1 Low-rank Adaptation Based on Random Vectors (RVLoRA)
Recent studies have demonstrated the remarkable effectiveness of using random weights in deep learning models. On the one hand, introducing random vectors can significantly enhance a model’s adaptive capability, providing an effective mechanism for adjusting the model’s behavior without the need for extensive retraining or parameter tuning [15, 16]. On the other hand, several advanced deep learning techniques, including those focused on feature learning and transformations, rely heavily on affine transformations of learned features. Freezing these randomly initialized parameters has been shown to improve the network’s ability to capture and generalize these affine transformations more effectively [25]. Additionally, in monocular depth estimation tasks, an inherent challenge is scale ambiguity, as depth information can vary in different contexts. Introducing random vectors specifically for scale adjustment can greatly improve the model’s flexibility, enabling it to better adapt to varying scales during inference. Inspired by these insights, we propose a novel and parameter-efficient fine-tuning method: low-rank adaptation based on random vectors (RVLoRA).
@@
where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$, $\Lambda_{b}\in\mathbb{R}^{m\times m}$, $\Lambda_{a}\in\mathbb{R}^{r\times r}$, with rank $r\ll\min(m,n)$. In this method, the low-rank matrices $B$ and $A$ are trainable, while the scaling vectors $b\in\mathbb{R}^{m\times 1}$ and $a\in\mathbb{R}^{r\times 1}$ are randomly initialized and remain frozen during training. These vectors are formally represented as diagonal matrices $\Lambda_{b}$ and $\Lambda_{a}$. This approach effectively performs random scaling on the rows of the low-rank matrices $B$ and $A$, thereby enhancing the model’s adaptability across layers.
@@ III-B2 Residual Block Based on Depthwise Separable Convolution (Res-DSC)
Transformers offer advantages like dynamic attention, global context awareness, and strong generalization capabilities. However, they often struggle to extract fine details and local features. On the other hand, convolutions excel at capturing local features and come with inherent benefits such as shift, scale, and distortion invariance, though they are limited in their ability to perceive global context. Therefore, integrating CNNs with Transformers can create a synergistic effect, significantly enhancing the overall model performance.
Inspired by this, we designed a residual block based on depthwise separable convolution (Res-DSC), as illustrated in Fig. 4. This structure consists of three convolutional layers: the first layer is a 1$\times$1 convolution to reduce the number of channels; the second layer is a 3$\times$3 depthwise separable convolution which has fewer parameters, leading to not only faster computation but also a more streamlined model compared to traditional convolution; the final layer is a 1$\times$1 convolution to restore the number of channels. To reduce the model’s parameters while ensuring the accuracy of depth estimation, we only add Res-DSC after the 3rd, 6th, 9th, and 12th transformer blocks. In Fig. 2, only the Res-DSC blocks after the 9th and 12th transformer blocks are shown, while the others are omitted for brevity.
@@ III-B3 Decoder
The decoder part adopts the same structure as Depth Anything Model [12]. The features from the last four layers of the encoder are fed into the decoder. Through the frozen Reassemble &\&& Fusion module in the decoder, features at different resolutions are obtained and then aggregated. These aggregated features are then fed into the trainable Depth-Head module for the final dense prediction, resulting in depth estimation outputs at four scales.
@@ III-C Total Training Framework
Figure 5: The visual results of the comparative experiments. It is evident that our method performs better in capturing details such as object edges. In particular, the regions highlighted by the black boxes demonstrate a significant advantage of our approach.Note:Since the ground truth in this dataset is captured using structured light encoding, comparing it with the depth map visually is not intuitive. As with other papers [7, 6, 13], we do not present the ground truth for comparison.
@@
SfMLearner [4] CVPR’17 0.086$\pm$0.037 1.021$\pm$1.100 7.553$\pm$5.110 0.121$\pm$0.060 0.925$\pm$0.081 31.60 31.60
Monodepth2 [5] ICCV’19 0.066$\pm$0.025 0.577$\pm$0.482 5.781$\pm$3.174 0.093$\pm$0.036 0.961$\pm$0.043 14.84 14.84
Endo-SfMLearner [20] MedIA’21 0.068$\pm$0.029 0.679$\pm$0.709 6.227$\pm$4.137 0.098$\pm$0.047 0.955$\pm$0.054 14.84 14.84
AF-SfMLearner [7] MedIA’22 0.060$\pm$0.024 0.443$\pm$0.347 4.964$\pm$2.348 0.082$\pm$0.032 0.973$\pm$0.037 14.84 14.84
IID-SfMLearner [6] JBHI’24 0.058$\pm$0.026 0.435$\pm$0.375 4.820$\pm$2.583 0.080$\pm$0.035 0.969$\pm$0.043 14.84 14.84
Depth Anything [12] CVPR’24 0.086$\pm$0.039 0.927$\pm$0.875 6.957$\pm$3.817 0.112$\pm$0.049 0.929$\pm$0.089 97.50 97.50
EndoDAC [13] MICCAI’24 0.051$\pm$0.022 0.355$\pm$0.334 4.442$\pm$2.509 0.073$\pm$0.032 0.979$\pm$0.032 99.09 1.66
Ours - 0.048$\pm$0.021 0.315$\pm$0.297 4.172$\pm$2.261 0.068$\pm$0.030 0.982$\pm$0.027 98.80 1.38
@@ IV-B Dataset and Metrics
We conducted our experiments on the SCARED dataset, which was collected using the da Vinci Xi endoscope during the dissection of fresh pig abdomens [30]. High-quality depth ground truth was obtained using structured light encoding during the acquisition process. Following the splitting strategy in prior works [6, 7], we divided the SCARED dataset into 15,351 frames for training, 1,705 frames for validation, and 551 frames for testing.
We adopted commonly used evaluation metrics such as Abs Rel, Sq Rel, RMSE, RMSE log, and Threshold $\delta$, with their calculation formulas shown in Table I. Similar to previous methods [6, 7], during the evaluation, we scale the predicted depth map via the median scaling, which can be expressed as:
@@
where $d$ and $d^{*}$ represents the predicted depth and the ground truth, respectively. The upper limit for scaling the depth map is set to 150 millimeters, which effectively covers almost all pixels in the SCARED dataset.
Figure 6: Qualitative comparison of applying the model trained on the SCARED dataset to the Hamlyn dataset for validation without any fine-tuning. Our method not only avoids erroneous estimates but also delivers more precise details. The regions highlighted by the black boxes illustrate a notable advantage of our approach, further demonstrating its strong generalization ability.
@@ IV-C Comparison Experiments
We compared our method with seven competitive unsupervised depth estimation methods: SfmLearner [4], Monodepth2 [5], Endo-SfmLearner [20], AF-SfmLearner [7], IID-SfmLearner [6], Depth Anything [12], and EndoDAC [13]. Among these, EndoDAC is currently the state-of-the-art method on the SCARED dataset. The first three methods are reproduced from scratch following the code provided by the authors, and the last four methods are tested using the optimal model provided by the authors. Table II presents the results of the comparative experiments.
@@
Fig. 5 presents some representative results for the compared methods. It can be observed that the depth maps estimated by our method are smoother and more continuous, with better performance at object boundaries, indicating that our method exhibits superior performance in capturing image detail information.
@@
TABLE IV: Experimental results on the Hamlyn dataset. The best results are presented in bold.
[…]
AF-SfMLearner [7] 0.130 2.134 11.306 0.165 0.843
IID-SfMLearner [6] 0.120 1.833 10.362 0.156 0.867
EndoDAC [13] 0.103 1.422 9.082 0.130 0.906
Ours 0.094 1.215 8.318 0.120 0.924
Like other methods [7], to further test the generalization ability of the proposed method, we directly applied the model trained on SCARED to the Hamlyn dataset for validation without any fine-tuning. The Hamlyn dataset is a laparoscopic/endoscopic video dataset created by the Hamlyn Centre at Imperial College London, containing rich laparoscopic and endoscopic video data [31].
The test results on Hamlyn dataset are shown in the Table IV. We only compare with a few of the most competitive methods. It can be seen that our method demonstrates strong generalization ability and continues to achieve state-of-the-art performance on this dataset. The visual results, as shown in the Fig. 6, clearly illustrate that the depth maps generated by our method exhibit greater continuity and superior performance in capturing object boundaries and texture details. This experiment provides strong evidence of the robust generalization capability of our approach.
Figure 7: Visualization of pose estimation results. The red line represents the ground truth, the green line represents the result from EndoDAC, and the blue line represents our method’s result. It can be observed that our results exhibit less trajectory drift and align more closely with the ground truth.
@@ IV-D Ablation Study
To further demonstrate the effectiveness of the proposed modules, ablation experiments were conducted and the results are shown in Table V. The first row in the table represents the results obtained by directly testing the Depth Anything Model without any fine-tuning. The subsequent rows show the results of progressively adding the proposed two modules. The experimental results indicate that both modules effectively improve the depth estimation performance.
In our method, the Res-DSC modules are evenly distributed across the 12 transformer modules of the encoder. For instance, when using three Res-DSC modules, they are placed at the 4th, 8th, and 12th layers. To determine the optimal number of Res-DSC modules, we conducted a series of ablation experiments. The results, shown in Table VI, indicate that using four Res-DSC modules achieves the best performance while maintaining a relatively low parameter count.
TABLE V: The ablation study results.
@@
TABLE VI: The ablation study on the number of Res-DSC modules
@@
TABLE VII: The results of pose estimation.
Method ATE (Seq 1$\downarrow$ ATE (Seq 2$\downarrow$ ATE (Seq 3$\downarrow$
EndoDAC [13] 0.0521 0.0390 0.0779
Ours 0.0505 0.0371 0.0771